A straight line passes through the point (2, –1, –1). It is parallel to the plane
4x + y + z + 2 = 0 and is perpendicular to the line


       x/1=y/(-2) =(z-5)/1                     Find the equations of the straight line ?


Grade:

1 Answers

Amit Saxena
14 years ago

Dear Sagar,

Let direction cosines of straight line be l, m, n

                              then :    4l + m + n = 0

                                          and for that of the given equation is

                                       :    l – 2m + n = 0

 thus ,

             l/3 = m/-3= n/-9  →  l/-1=m/-3=n/3.

askIITians.com provides online iit jee courses and IIT JEE Test Series with IITians. Click here to get free online test series and check your status timely or you can join us as our registered user for getting best iit jee study material or iit jee test series.

Click here to find IIT JEE Packages or course details with askIITians.

for further details visit the folowing links:

http://www.askiitians.com//iit-jee-online-coaching.aspx

http://www.askiitians.com/iit-jee-test-series.aspx

http://www.askiitians.com/iit-study-material/

Askiitians provide you all the best possible information, feel free to ask.

all the best.

thanks and regards.

Akhilesh Shukla